OpenStudy (anonymous):

Which of the following best defines 2 to the power of 2 over 3? Cube root of 4 Square root of 4 Cube root of 2 Square root of 2

OpenStudy (astrophysics):

\[\large x ^{\frac{ m }{ n }} = \sqrt[n]{x ^{m}}\]

OpenStudy (astrophysics):

\[\huge x ^{\frac{ 2 }{ 3 }}\] You have this, what will it be giving what I showed you?